# FROST-BRDF: A Fast and Robust Optimal Sampling Technique for BRDF Acquisition

### v1

@@ III-A Problem formulation
Due to the high-dimensional nature of measured BRDFs, an exhaustive search through all the sensing matrices within $\Psi$ is not possible for solving (6). For instance, the set of sub-sampling operators, $\Psi$, considering $20$ samples from a BRDF with a resolution of $90\times 90\times 180$, contains (${1458000\choose 20}\approx 10^{104}$ possible candidates for the optimal sub-sampling operator.
@@ III-B Solution via FROST
respectively, where .$\|.\|_{r0}$ is the $\ell_{0}$ norm defined over rows, i.e. the number of nonzero rows. Equations (8) and (9) are classical examples of sparse representation under the MMV model, where we assume that the same support is shared among all signals. The variable $\mathbf{Z}$ is a replacement for the training set $\mathbf{T}$, which is the unknown according to the analysis formulation in (7). However, since we are only interested in optimal sample locations for the training set, i.e. the rows of $\mathbf{T}$, we do not need to recover the matrix $\mathbf{Z}$, but rather the row support of $\mathbf{Z}$. We denote the row support of $\mathbf{Z}$ by $\lambda$. The set $\lambda$ defines the optimal sample locations for the training set $\mathbf{T}$. Note that while the sub-sampling operator $\mathbf{\Phi}$ does not explicitly appear in (8) and (9), it is implicitly defined using the support set $\lambda$. In other words, one can construct a sub-sampling operator, from an identity matrix as described above, using the support set $\lambda$. This demonstrates the equivalence of (5) and (6) with (8) and (9), respectively. The index set $\lambda$ can be used in, e.g., a gonioreflectometer to obtain a set of BRDF samples for recovering the full BRDF by solving (1). More details on the reconstruction algorithm and the choice of the dictionary $\mathbf{D}$ will be presented in Section IV.
@@
3:$l\leftarrow 1$
4:$\lambda^{(0)}\leftarrow\varnothing$triangleright The support set at iteration 0
5:$\mathbf{R}^{(0)}\leftarrow\mathbf{S}$triangleright The residual at iteration 0
@@
8:   $\lambda^{(l)}\leftarrow\lambda^{(l-1)}\cup\{j\}$triangleright Support set update
9:   $\mathbf{R}^{(l)}\leftarrow\mathbf{S}-(\mathbf{D}^{-1}_{\lambda^{(l)}})(\mathbf{D}^{-1}_{\lambda^{(l)}})^{\dagger}\mathbf{S}$triangleright Residual update
10:   $l\leftarrow l+1$
@@ IV-B Reconstruction algorithm
Figure 11: Visual comparison results for FROST-BRDF and the method of Xu et al. [32] using $20$ sample directions. From left to right, the materials are acrylic-felt-orange-rgb, satin-rosaline-rgb, and irid-flake-paint1-rgb, all from the RGL-EPFL data set. We used the MERL data set for the training of both methods. Rendering SNR is reported below each image.
@@ V Results and Discussions
We randomly split the dataset into 131 materials for training, and 20 materials for testing as our base experiment setting. Out of 131 materials included in the training set, 86 and 45 materials were collected from the MERL and RGL-EPFL datasets, respectively. We evaluated our results with a 10-fold cross-validation each with an individual random training - test set split across the datasets. In this setup, the training data matrix $\mathbf{T}$, introduced in Section III-A, has $n=90\times 90\times 180$ rows, and $t=131\times 3=393$ columns, since each BRDF has three color channels. Note that the total number of valid BRDF values is $1,111,432$. Therefore, in practice, the size of the matrices $\mathbf{T}$ and $\mathbf{D}$ are $1,111,432\times 393$, which amounts to $1.63$GB of storage, assuming a $32$-bit floating point storage.
@@
Our method was implemented in MATLAB and the source code for FROST-BRDF accompanies this manuscript. After the reconstruction of a full BRDF using FROST from optimal samples, we render the result using PBRT [68], followed by tone mapping the image with a Gamma function, where $\gamma=2.0$. The HDR environment map used here is Grace Cathedral. The absolute error images were normalized with respect to the maximum absolute pixel error for our method and those we compare to.
@@ V-C Visual quality and number of samples
Figure 8, shows four examples of the reconstructed BRDFs with $m=20$ samples. The results from Nielsen et al. [13] appear in the first row, and those from our FROST-BRDF method are in the second row. The glossy materials, such as specular-blue-phenolic and gold-metallic-paint2, and the diffuse materials, such as green-latex, can be reconstructed faithfully with our FROST-BRDF compared to the baseline method. Even in the layered material, cm-white-rgb, that is known to be challenging to model, our method achieves a rendering SNR of $32.16$dB.

### v2

@@ III-A Problem formulation
Due to the high-dimensional nature of measured BRDFs, an exhaustive search through all the sensing matrices within $\Psi$ is not possible for solving (6). For instance, the set of sub-sampling operators, $\Psi$, considering $20$ samples from a BRDF with a resolution of $90\times 90\times 180$, contains ${1458000\choose 20}\approx 10^{104}$ possible candidates for the optimal sub-sampling operator.
@@ III-B Solution via FROST
respectively, where $\|.\|_{r0}$ is the $\ell_{0}$ norm defined over rows, i.e. the number of nonzero rows. Equations (8) and (9) are classical examples of sparse representation under the MMV model, where we assume that the same support is shared among all signals. The variable $\mathbf{Z}$ is a replacement for the training set $\mathbf{T}$, which is the unknown according to the analysis formulation in (7). However, since we are only interested in optimal sample locations for the training set, i.e. the rows of $\mathbf{T}$, we do not need to recover the matrix $\mathbf{Z}$, but rather the row support of $\mathbf{Z}$. We denote the row support of $\mathbf{Z}$ by $\lambda$. The set $\lambda$ defines the optimal sample locations for the training set $\mathbf{T}$. Note that while the sub-sampling operator $\mathbf{\Phi}$ does not explicitly appear in (8) and (9), it is implicitly defined using the support set $\lambda$. In other words, one can construct a sub-sampling operator, from an identity matrix as described above, using the support set $\lambda$. This demonstrates the equivalence of (5) and (6) with (8) and (9), respectively. The index set $\lambda$ can be used in, e.g., a gonioreflectometer to obtain a set of BRDF samples for recovering the full BRDF by solving (1). More details on the reconstruction algorithm and the choice of the dictionary $\mathbf{D}$ will be presented in Section IV.
@@
[…]
@@
[…]
@@ IV-B Reconstruction algorithm
Figure 11: Visual comparison results for FROST-BRDF and the method of Xu et al. [32] using 2020 sample directions. From left to right, the materials are acrylic-felt-orange-rgb, satin-rosaline-rgb, and irid-flake-paint1-rgb, all from the RGL-EPFL data set. We used the MERL data set for the training of both methods. Rendering SNR is reported below each image.
@@ V Results and Discussions
We randomly split the dataset into 131 materials for training, and 20 materials for testing as our base experiment setting. Out of 131 materials included in the training set, 86 and 45 materials were collected from the MERL and RGL-EPFL datasets, respectively. We evaluated our results with a 10-fold cross-validation each with an individual random training - test set split across the datasets. In this setup, the training data matrix $\mathbf{T}$, introduced in Section III-A, has $n=90\times 90\times 180$ rows, and $t=131\times 3=393$ columns, since each BRDF has three color channels. Note that the total number of valid BRDF values is $1,111,432$. Therefore, in practice, the size of the matrices $\mathbf{T}$ and $\mathbf{D}$ are $1,111,432\times 393$, which amounts to 1.631.63GB of storage, assuming a $32$-bit floating point storage.
@@
Our method was implemented in MATLAB and the source code for FROST-BRDF accompanies this paper [^1] https://github.com/emiandji/frost_brdf. After the reconstruction of a full BRDF using FROST from optimal samples, we render the result using PBRT [68], followed by tone mapping the image with a Gamma function, where $\gamma=2.0$. The HDR environment map used here is Grace Cathedral. The absolute error images were normalized with respect to the maximum absolute pixel error for our method and those we compare to.
@@ V-C Visual quality and number of samples
Figure 8, shows four examples of the reconstructed BRDFs with $m=20$ samples. The results from Nielsen et al. [13] appear in the first row, and those from our FROST-BRDF method are in the second row. The glossy materials, such as specular-blue-phenolic and gold-metallic-paint2, and the diffuse materials, such as green-latex, can be reconstructed faithfully with our FROST-BRDF compared to the baseline method. Even in the layered material, cm-white-rgb, that is known to be challenging to model, our method achieves a rendering SNR of 32.1632.16dB.
@@ VII Conclusions
VIII Acknowledgement
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 956585 (PRIME).